• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 23:57:10 on January 31, 1999

    Entry number 11941

    keyword=Beam Time Accounting
    Hall A

    1/31/1999 1600-2400
    Experiment #: E95-001
    Shift Leader: J. P. Chen
    Shift Workers: A. Deur, M. Finn, J. Martin

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    30 40 1 40 1 45 1 1 6.58333333333333
    Beam Available
    but Not in Use
    0 0 0 20 0 15 0 0 0.583333333333333
    Beam Not Available
    or Unacceptable
    30 20 0 0 0 0 0 0 0.833333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 40 1 45 1 1 7.41666666666667
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 20 0 15 0 00.583333333333333
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 20 0 15 0 0 0.583333333333333

    Comments: smooth running 16:00-16:30 Q2=0.1 16:30 system check and NMR measurement pol=29.9% also did pedestal run. 17:05 beam back, raster check and slow ramping beam 17:25 data taking again (Q2=0.1) fast feed back was off and (after MCC fixed problem) tuned back on at 17:55. 19:00 Harp Scan, and then e-P, but e-P had problem, did not get any event. 19:30 raster checking, ramping beam 19:40 resume regular data taking (Q2=0.1). 21:10 Get ready for Moller run/check e-p one more time. 21:25 -2400 Moller measuremnet took the rest of the shift.

    Data Entry by: J. P. Chen